Q#1 For any sets 𝐴, 𝐵 and 𝐶, if 𝐴 ⊆ 𝐵, then 𝐴 ∩ 𝐶 ⊆ 𝐵 ∩ 𝐶.

Q#2 For any sets 𝐴, 𝐵 and 𝐶,

(𝐴 × 𝐶) ∩ (𝐵 × 𝐷) = (𝐴 ∩ 𝐵) × (𝐶 ∩ 𝐷).

Q#3 Given sets 𝐴, 𝐵 and 𝐶, prove that

𝐴 × (𝐵 ∩ 𝐶) = (𝐴 × 𝐵) ∩ (𝐴 × 𝐶)

Q#4 Prove that If 𝐴 and 𝐵 are sets, then 𝒫(𝐴)⋃𝒫(𝐵) ⊆ 𝒫(𝐴⋃𝐵).

Q#5 Suppose 𝐴 and 𝐵 are sets.

If 𝒫(𝐴) ⊆ 𝒫(𝐵),𝑡ℎ𝑒𝑛 𝐴 ⊆ 𝐵.


1
Expert's answer
2021-10-21T13:26:15-0400

Q#1.

Let "x\in A\cap C"

Case 1,

Since "A\sube B, x\in A\implies x\in B"

"x\in B\implies x\in B\cap C"

Case 2,

"x\in C"

"x\in C\implies x\in B\cap C" since "x\in B"

Thus, in either cases, "x\in A\cap C\implies x\in B\cap C."

So,

"\forall x(x\in A\cap C\implies x\in B\cap C)"

We now have that,

"A\cap C\sube B\cap C" "\blacksquare."


Q#2.

Let "(x,y)\in(A\cap B)\times(C\cap D)"

"\iff x\in (A\cap B)\times(C\cap D)\times y"

"\iff x\in A\cap B\land y\in C\cap D"

"\iff x\in A\land x\in B\land y\in C\land y\in D"

"\iff x(A\times C)y\land x(B\times D)y"

"\iff x(A\times C)\cap(B\times D)y"

Therefore,

"(A\cap B)\times(C\cap D)=(A\times C )\cap (B\times D)" as required "\blacksquare" .


Q#3.

To begin with, we will first show that "A\times (B\cap C)\sube (A\times B)\cap(A\times C )". Suppose "(x,y)\in A\times(B\cap C)". This means that, "x\in A" and "y\in B\cap C" by definition of Cartesian product. By definition of intersection, it follows that "y\in B" and "y\in C".Since "x\in A" and "y\in B", it follows that "(x,y)\in A\times B" by definition of Cartesian product. Also, since "x\in A" and "y\in C" it follows that "(x,y)\in A\times C" by definition of "\times". Now, we have "(x,y)\in A\times B" and "(x,y)\in A\times C" and so, "(x,y)\in (A\times B)\cap(A\times C)". Having shown that "(x,y)\in A\times(B\cap C)\implies (x,y)\in (A\times B)\cap(A\times C)", we have, "A\times(B\cap C)\sube(A\times B)\cap(A\times C)".

Next, we show that "(A\times B)\cap(A\times C)\sube A \times(B\cap C)",

Suppose "(x,y)\in (A\times B)\cap(A\times C)". By definition of intersection, it means that, "(x,y)\in A\times B" and "(x,y)\in A\times C". By definition of Cartesian product, "(x,y)\in A\times B" means "x\in A" and "y\in B" also, "(x,y)\in A\times C" means "x\in A" and "y\in C". We have that, "y\in B" and "y\in C" so, "y\in B\cap C" by definition of intersection. Thus, we have deduced that "x\in A" and "y\in B\cap C", so "(x,y)\in A\times (B\cap C)". Hence,

"(A\times B)\cap(A\times C)\sube A\times (B\cap C)".

In summary, we have shown that "A\times (B\cap C)\sube (A\times B)\cap(A\times C)" and "(A\times B)\cap(A\times C)\sube A\times (B\cap C)". It follows that, "A\times(B\cap C)=(A\times B)\cap(A\times C)" as required"\blacksquare."


Q#4.

Suppose "x\in \mathcal{p}(A)\cup \mathcal{p}(B)",

Then, "x\in\mathcal{p}(A) \space or\space x\isin\mathcal{p}(B)"

Hence,

"x\sube A\space or\space x\sube B"

We need to show that "x\in\mathcal{p}( A\cup B)," that is, we need to show that "x\sube A\cup B"

Suppose "y\in x" then, "y\in A\space or\space y\in B" since "x\sube A \space or \space x\sube B"

Thus, "y\in x\implies y\in (A\cup B)" and so, "x\sube (A\cup B)".

Hence,

"x\in\mathcal {p}(A\cup B)"

"\therefore x\in \mathcal{p}(A)\cup \mathcal{p}(B)\implies x\in\mathcal {p}(A\cup B)"

and so,

"\mathcal{p}(A)\cup\mathcal{p}(B)\sube \mathcal{p}(A\cup B)" as required "\blacksquare" .


Q#5.

Suppose that "A" and "B" are sets and "\mathcal p(A)\sube\mathcal {p})(B)"

By definition of a power set, "A\sube \mathcal {p}(A)".Since "A\in \mathcal{p}(A)" and "\mathcal{p}(A)\sube \mathcal{p}(B)", we know that "A\in \mathcal{p}(B)" by definition of subset.

So, by definition of power set, "A\sube B" "\blacksquare".